estimate how much water(2) composition is in the air(1) during the night (T=20C, Psat water=60.09 kPa) and during the day (T= 35C, Psat water= 72.5 kPa). also calculate how much air is dissolved in water at 20C if henry's constant is 7,250,000 kPa. assume that the pressure is 101.33 kPa.
Answer #1

Henry's constant H = 7250,000 kPa

Total pressure P = 101.33 kPa

During the night condition:

T = 20 C, Psat = 60.09 kPa

During the day

T = 35 C and Psat = 72.5 kPa

Sol:

for night,

Composition of water in air , y= Psat/P = 60.09/101.33 = 0.593

for day,

Composition of water in air, y = Psat/P = 72.5/101.33 = 0.715

how much air is dissolved into water at 20C,

by henry law,

P = H*x

x = P/H = 101.33/7250000 = 1.39E-5

Air is dissolved into air = 1.39E-5